OpenStudy (anonymous):

Which is 243 written as a power of a number?

OpenStudy (jojo4eva):

3^5

OpenStudy (jojo4eva):

3*3=9*3=27*3=81*3=243

OpenStudy (anonymous):

thanks got an 80